Circular No. 8709

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006cb, 2006cc, 2006cd, 2006ce Four supernovae have been found in unfiltered images: 2006cb, 2006cc,

and 2006cd by LOSS on KAIT images by N. J. Ponticello, H. Khandrika, D. R. Madison, and W. Li (cf. IAUC 8707); 2006cd independently by J. Newton, T. Crowley, and T. Puckett with a 0.40-m reflector at Portal, AZ, in the course of the Puckett Observatory Supernova Search (the tabulated data below for 2006cd is that from Puckett); and 2006ce by L. A. G. Monard, Pretoria, South Africa (0.30-m reflector; cf. IAUC 8699).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006cb May 5.43 14^{h}16^{m}31.^{s}80 +39^{o}35′15′′.9 17.5 0′′.6 E, 5′′.7 S
2006cc May 6.45 16 09 56.47 +43 07 35.9 18.3 11′′.1 E, 9′′.4 S
2006cd May 8.27 16 05 22.06 +17 45 05.5 18.3 5′′.6 W, 11′′.4 S
2006ce May 10.14 2 22 54.63 −21 14 29.4 12.4 136′′ W, 28′′ S

Additional magnitudes for 2006cb in NGC 5541: Feb. 23.48 UT, [19.5; Apr. 19.39, 18.6. Additional magnitudes for 2006cc in UGC 10244: Feb. 13.48, [19.5; Apr. 26.45, [18.5; May 7.43, 18.1. SN 2006cc is a type-Ia supernova, near maximum on May 8 (cf. CBET 506). Additional magnitudes for 2006cd in IC 1179: 2005 Sept. 9, [19.7 (Puckett); 2006 Feb. 23.45, [19.5 (KAIT); May 8.44, 18.2 (KAIT); May 9.23, 18.2 (A. Sehgal, 0.50-m reflector, Osoyoos, BC); 9.45, 18.1 (KAIT). Li gives the offset for 2006cd as 3′′.0 west and 11′′.2 south of the center of IC 1179, the southwest component of Arp 272 (the northeast component being NGC 6050). Additional magnitudes for 2006ce in NGC 908: Apr. 1.713, R > 16.5 (not detected); May 11.151, B = 12.9, V = 12.55, R = 12.4. Also, nothing is visible at the location of 2006ce on the Digitized Sky Survey (limiting red mag 20.5) or on Monard’s images taken prior to Apr. 1 (limiting mag ∼ 18).

COMET 73P/SCHWASSMANN-WACHMANN As comet 73P has neared the earth, increasingly large numbers of

fragments (some apparently very short-lived) have been observed. The Minor Planet Center has continued to assign letter designations to those components observed on three or more nights with precise astrometry: ‘AQ’ and ‘AR’ announced on MPEC 2006-H37, ‘AS’ and ‘AT’ on MPEC 2006-H48, ‘AU’–‘BM’ on MPEC 2006-H61, and ‘BN’–‘BP’ on MPEC 2006-J31.

Visual total-magnitude estimates of component ‘B’ by J. J. Gonzalez, Leon, Spain: Apr. 28.91 UT, 7.6 (25×100 binoculars); May 1.02, 7.0 (7×50 binoculars); 5.12, 6.8 (0^{o}.6 tail in p.a. 235^{o}); 11.14, 5.2 (naked eye).

2006 May 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green